OpenStudy (anonymous):

I need help with solving this problem. The instructions are: Given f(x) = 2x + 5 and g(x) = x2 + 2, compute the following. Simplify your answers. Here's the problem. There are several parts to it to solve. a) g(–5) (b) f(a + h) (c) g(a) Any assistance would be GREATLY APPRECIATED!!

OpenStudy (anonymous):

for a all you do is supplement the -5 in for x

OpenStudy (anonymous):

Im not 100 % sure though

OpenStudy (agent0smith):

a) assuming it's x^2 not x*2, substute in -5 for x \[g(-5) = (-5)^2 + 2\] b) substitute (a+h) in for x \[ f(a+h) = 2(a+h)+ 5\] c) substitute in a for x \[ g(a) = a^2 + 2\]

OpenStudy (anonymous):

Just substitute the value for x in each function: g(-5)= (-5)^2 + 2 = ? f(a+h) = 2(a+h)+5 = ? g(a) = a^2 + 2 = ?

OpenStudy (agent0smith):

And then simplify.

OpenStudy (anonymous):

All of the above answers are correct, but it is important to realize why. f(x) notation means that you are looking at a function of x. You have been dealing with these a lot so far, but in the form y=mx+b. The value inside the parenthesis is the value of the variable which the function is in terms of. so if you had f(z) and you are asked to solve f(2) you would plug in 2 for all values of z. This notation is extremely important to understand as you will see it any form of math you followup with.

OpenStudy (anonymous):

Thank you to all who responded to my dilemma. I had tried to figure that problem out for three days prior to posting it on here with no success. I hope all the members of OpenStudy have a Merry Christmas and a Happy and Prosperous New Year in 2013!